the sum of three numbers is 64. the second number is 3 more than the first. the third number is 11 less than twice the first. find the numbers.
Answer by oberobic(2304)

Three numbers can be defined as x, y, and z.
x + y + z = 64
y = x+3
z = 2x - 11
.
substitute for y and z
x + (x+3) + (2x-11) = 64
4x -8 = 64
4x = 72
x = 18
.
y = x+3 = 21
z = 2x -11 = 2(18) -11 = 36-11 = 25
.
x + y + z = 18 + 21 + 25 = 64
Correct.
.
Answer: The three numbers are 18, 21, and 25.
.
Done.